OpenStudy (megannicole51):

Find the area lying outside r=6cos(theta) and inside r=3+3cos(theta)

zepdrix (zepdrix):

So in polar we'll use:\[\Large A \quad=\quad \frac{1}{2} \int\limits r^2\;d \theta\] So ummm, we need an r. Have you tried graphing these functions? :) https://www.desmos.com/calculator/jgur6caz6i So we're interesting in the area outside of that red circle, but inside of the blue, yes?

zepdrix (zepdrix):

So that radial length is given by the blue function minus the red. \[\Large r=3+3\cos \theta-6\cos \theta\]

OpenStudy (anonymous):

@zepdrix, for the area between two curves we need an outer \(r\) and an inner \(r\). In this case, \(r=6\cos\theta\) is the inner and \(r=3+3\cos\theta\) is the outer. Then, the area would be \[\frac{1}{2}\int_\alpha^\beta(r_\text{outer}^2-r_\text{inner}^2)~d\theta\]

OpenStudy (megannicole51):

i like your demo @zepdrix thats super cool

OpenStudy (megannicole51):

r=3-cos(theta)?

zepdrix (zepdrix):

Oh did I make a boo boo there? :( lol sorry, just woke up. the graph megan? :O ikr? cool little tool

OpenStudy (anonymous):

just a slight one : )

OpenStudy (megannicole51):

(1/2) integral of (6cos(theta))^2 - (3+3cos(theta))^2 ????

zepdrix (zepdrix):

Hmm I think you might have those backwards. The outer function is r=3+3cos theta. We would want to subtract the smaller from that.

OpenStudy (megannicole51):

the outer function is 6cos(theta)

zepdrix (zepdrix):

It is? D: See how the blue is OUTSIDE of the red?

zepdrix (zepdrix):